OpenStudy (anonymous):

solve for y: 3x+2y=16

OpenStudy (lalaly):

y=16-3x/2

OpenStudy (angela210793):

2y=16-3x y=8-3x/2

OpenStudy (anonymous):

so which on is it..

OpenStudy (angela210793):

they are the same each of them is right

OpenStudy (lalaly):

(16-3x) /2 and 8 - (3x/2)